# Quintessential early dark energy

### v2

@@ 4.2 Background dynamics
Figure 3: Numerically evolved MSE potential (4.1) for $V_{0}=5\times 10^{10}\rho_{\rm c0}$, $n=6$, $\phi_{0}=0.7$ and $\Omega_{\mathrm{m}0}=0.3$. Initial conditions are $\phi_{i}=\phi_{0}-0.1\phi_{0}$ and $\phi_{i}^{\prime}=0.1$.
@@
Figure 4: From Eq. (5.10) and for the MSE potential (4.1) $V_{0}=5\times 10^{10}\rho_{\rm c0}$, $n=6$, $\phi_{0}=0.7$ and $\Omega_{\mathrm{m}0}=0.3$. Initial conditions are $\phi_{i}=\phi_{0}-0.1\phi_{0}$ and $\phi_{i}^{\prime}=0.1$.
@@
As we have already discussed that for $\phi/M_{\textrm{Pl}}\ll\phi_{0}^{1/n}$, i.e., at higher redshifts (early times), the MSE potential (4.1) reduces to the SE potential (3.1). So, the cosmological dynamics of the scalar field will be similar to the SE potential at higher redshifts or early times. Also, when $\phi/M_{\textrm{Pl}}\sim\phi_{0}^{1/n}$ the MSE potential is still very steep which can be understood from Fig. 3 which shows the numerical evolution of the MSE potential for $n=6$ and $\phi_{0}=0.7$. For the chosen $\phi_{0}$ we have $\phi_{0}^{1/n}\approx 0.94$ where the potential is very steep. Note that the potential has been shown in log scale in the Fig. 3. So, around $\phi_{0}$ we will have the effect of the steep potential which can lead to multiple frozen epochs in the scalar field dynamics as discussed in the previous section (Sec. 3) and has been shown in the upper figure of Fig. 1. For $\phi/M_{\textrm{Pl}}>\phi_{0}^{1/n}$, from Eq. (4.4), we see that
$V(\phi)\approx V_{0}\mathrm{e}^{-\gamma}\mathrm{e}^{\gamma\phi_{0}M_{\textrm{Pl}}^{n}/\phi^{n}}\,.$ (4.9)
@@
The above equation is very crucial to understand the late time dynamics of the scalar field. The MSE potential (4.1) behaves like the potential (4.9) before becoming a flat potential i.e., during the recent past. The potential (4.9) is a generalization of the potential $V(\phi)\sim\mathrm{e}^{M_{\textrm{Pl}}/\phi}$ which gives rise to tracker solution [8, 9]. So, the potential (4.9) would also give rise to a tracker solution. In tracker behaviour the scalar field dynamics does not exactly follow the background energy density like in scaling behaviour but decays a little slower than the background. Because of this slower decay process the scalar field energy density eventually takes over the background unlike the scaling behaviour in which the scalar field energy density follows the background forever. This tracker behaviour in the MSE potential (4.1) brings a slight difference in the scalar field dynamics compared to the SE potential (3.1). Now when $\phi/M_{\textrm{Pl}}\gg\phi_{0}^{1/n}$ the MSE potential (4.1) behaves like a CC which makes the scalar field to exit from the tracker behaviour and the scalar field EoS $w_{\phi}$ becomes very close to $-1$. So, the MSE potential not only can give rise to viable cosmology but also give interesting dynamics along with relating two different energy scales and thereby we do not need to invoke any CC in the theory. The steep nature of the potential gives multiple frozen epochs in the evolution of the scalar field energy density which can be very useful to achieve viable EDE solutions which we discuss in the next section (Sec. 5). In Fig. 4 we have shown the cosmological evolution of different cosmological parameters for the MSE potential (4.1).In the top right figure of Fig. 4 we see that we have two frozen epochs in the evolution of the scalar field energy density which is followed by a tracker behaviour. The bottom figure of Fig. 4 shows that the scalar field EoS $w_{\phi}$ is very close to $-1$ at $z=0$. The top left figure of the Fig. 4 shows the evolution of the density parameters.
@@ 5 Early Dark Energy
$\theta_{\rm s}=\frac{r_{\rm s}(z_{\star})}{(1+z_{\star})D_{\rm A}(z_{\star})}\,,$ (5.1)
@@
$r_{\rm s}(z)$ $=$ $\int_{z}^{\infty}\frac{c_{s}(z)}{H(z)}{\rm d}z=\frac{1}{H_{\star}}\int_{z}^{\infty}\frac{c_{s}(z)}{H(z)/H_{\star}}{\rm d}z\,,~{}~{}~{}$ (5.2)
$D_{\rm A}(z)$ $=$ $\frac{c}{1+z}\int_{0}^{z_{\star}}\frac{dz}{H(z)}=\frac{c}{H_{0}(1+z)}\int_{0}^{z}\frac{dz}{E(z)}\,,~{}~{}~{}~{}$ (5.3)
@@
$c_{\rm s}(z)$ $=$ $\frac{c}{\sqrt{3(1+\frac{R_{\rm b}}{1+z})}}\,,$ (5.4)
$R_{\rm b}$ $=$ $\frac{3\omega_{\rm b}}{4\omega_{\gamma}}\,,$ (5.5)
$\omega_{\mathrm{r}}$ $=$ $\left(1+\frac{7}{8}N_{\rm eff}\left(\frac{4}{11}\right)^{4/3}\right)\omega_{\gamma}\,,$ (5.6)
@@
$z_{\star}$ $=$ $1048\left(1+0.00124\omega_{\rm b}^{-0.738}\right)\left(1+g_{1}\omega_{\mathrm{m}}^{g_{2}}\right)$ (5.7)
$g_{1}$ $=$ $\frac{0.0783\omega_{\rm b}^{-0.238}}{1+39.5\omega_{\rm b}^{0.763}}$ (5.8)
$g_{2}$ $=$ $\frac{0.560}{1+21.1\omega_{\rm b}^{1.81}}\;.$ (5.9)
@@
Table 1: Value of $H_{0}$ in the unit of ${\rm km}\;{\rm s}^{-1}{\rm Mpc}^{-1}$ for different $n$ by maintaining $\Omega_{\mathrm{m}0}\approx 0.31$, $f_{\rm EDE}\approx 0.07$, $z_{\star}\approx 1091$ and $z_{\rm eq}=3300$ and the redshift of drag epoch $z_{\rm d}=1060$. To calculate $H_{0}$ we have added a CC with SE (3.1), axionlike (5.12) and power law (5.13) potentials. Here $\lambda$ signifies $\mu$, $\phi_{0}$ and $f$.
@@
Using Eqs. (5.1), (5.2) and (5.3) we can have [11]
$H_{0}$ $=$ $\sqrt{3}H_{\star}\theta_{\rm s}\frac{\int_{0}^{z_{\star}}\frac{{\rm d}z}{E(z)}}{\int_{z_{\star}}^{\infty}\frac{E(z_{\star})}{E(z)}(1+R)^{-1/2}{\rm d}z}$ (5.10)
@@
$H_{\star}$ $=$ $\omega_{\mathrm{r}}^{1/2}(1+z_{\rm ls})^{2}\Big{(}\frac{\omega_{\mathrm{m}}}{\omega_{\mathrm{r}}}\times\frac{1}{1+z_{\star}}+1\Big{)}^{1/2}\times 100{\rm km~{}sec^{-1}Mpc^{-1}}\,.$ (5.11)
Eq. (5.10) gives us an estimate of $H_{0}$. We have listed values of $H_{0}$ for different parameter values for the SE (3.1) and MSE (4.1) potentials in Tab. 1 along with the values of $H_{0}$ for the axion-like and power law potentials given by
$V(\phi)$ $=$ $V_{0}\left(1-\cos\left(\frac{\phi}{f_{\rm pl}}\right)\right)^{n}\,,$ (5.12)
$V(\phi)$ $=$ $V_{0}\left(\frac{\phi}{f_{\rm pl}}\right)^{2n}\,,$ (5.13)
respectively by keeping the EDE density parameter at the redshift $z_{\rm c}$, $f_{\rm EDE}$ fixed at $0.07$ as small amount of $f_{\rm EDE}(<0.1)$ is allowed by the observation [40, 41, 38, 43, 24]. In the above equations $f_{\rm pl}=fM_{\textrm{Pl}}$ where $f$ is a constant. One should note that to estimate $H_{0}$ for using the Eq. (5.10) we are using the bounds from CMB observations on the parameters $\omega_{\mathrm{r}}$, $\omega_{\gamma}$, $\omega_{\mathrm{m}}$, $\omega_{\rm b}$ and $\theta_{\rm s}$ as mentioned above. So, this is just an estimation. The proper analysis should be done by doing observational data analysis by considering different cosmological observations. From Tab. 1 we see that the MSE potential (4.1) can improve the value of $H_{0}$ compared to $\Lambda$CDM but SE potential (3.1) with a CC improves the value even more. This happens because of the presence of tracker behaviour in the MSE potentials during the late time while we have scaling behaviour in the SE potential during the late time. Due to tracker behaviour the present value of $w_{\phi}$ can be slightly higher than $-1$ which can reduce the effect of EDE. While in SE potential, as we are adding a CC, the present value of $w_{\phi}$ is $-1$ which can have maximum effect of EDE for $w_{\phi}\geq 0$. It is also should be noted from Tab. 1 that the EDE in SE potential can give us better result in terms of the value of $H_{0}$ compared to the axion-like potential. To get the proper estimation of $H_{0}$ we do observational data analysis in the next section.
@@ 6 Growth of structures
$\ddot{\delta}+2H\dot{\delta}-4\pi G\bar{\rho}_{m}\delta=0\,,$ (6.2)
where, $\bar{\rho}_{\mathrm{m}}$ is the unperturbed (background) matter energy density and $\delta=(\rho_{\mathrm{m}}-\bar{\rho}_{\mathrm{m}})/\bar{\rho}_{\mathrm{m}}$ is the density contrast. The solution of Eq. (6.2) can be written as the linear superposition of the growing $D_{+}$ and decaying $D_{-}$ modes,
$\delta(t,\vec{k})=c_{+}D_{+}(t)\delta(\vec{k},0)+c_{-}D_{-}(t)\delta(\vec{k},0)\,,$ (6.3)
@@
$f=\frac{\mathrm{d}\ln D_{+}}{\mathrm{d}\ln a}\,.$ (6.4)
@@
$\Big{<}\delta(t,\vec{k})\delta(t,\vec{k}^{\prime})\Big{>}=\delta^{(3)}(\vec{k}+\vec{k}^{\prime})\mathcal{P}(t,k)\,,$ (6.5)
@@
$\mathcal{P}(t,k)$ $\propto$ $A_{\rm H}^{2}\Big{|}D_{+}(\tilde{a})\Big{|}^{2}T^{2}(k)\left(\frac{k}{H_{0}}\right)^{n_{\mathrm{s}}}\,,$ (6.6)
@@
$\sigma_{\rm R}^{2}=\int_{0}^{\infty}\mathrm{d}k\frac{k^{2}}{2\pi^{2}}\mathcal{P}(t,k)\big{|}W_{\rm win}(kR)\big{|}^{2}\,,$ (6.7)
@@
$\delta(\vec{x};R)=\int\delta(\vec{x}^{\prime})W_{\rm win}(\vec{x}-\vec{x}^{\prime};R)\mathrm{d}^{3}x^{\prime}\,.$ (6.8)
@@
$W_{\rm win}(kR)=\frac{3}{(kR)^{3}}\Big{(}\sin(kR)-kR\cos(kR)\Big{)}\,,$ (6.9)
The smoothing scale at which $\sigma_{\rm R}\sim 1$ is the scale where the linear perturbation theory breaks down and nonlinear effects become important. In this regard $R=8h^{-1}\rm Mpc$ is a relevant scale. From Planck 2018 results we have $\sigma_{8}=0.811\pm 006$ [3] at $z=0$. Using the best fit value of $\sigma_{8}$ from Planck 2018 results we fix the normalization factor $A_{\rm H}$ of the power spectrum  (6.6). $\sigma_{8}(z)$ can be represented by the growing mode $D_{+}(z)$
$\sigma_{8}(z)=\sigma_{8}(0)\frac{D_{+}(z)}{D_{+}(0)}\,.$ (6.10)
@@
Figure 5: Left: Power spectrum is shown for the MSE potential (4.1) with $n=6$ (solid blue) and $n=10$ (dotted green) at $z=0$ along with the power spectrum for the $\Lambda$CDM (dashed red) model. Right: Evolution of $f\sigma_{8}(z)$, for the similar cases as the left figure, has been shown. The brown dots are the observational data of $f\sigma_{8}$ along with their $1\sigma$ error bars [83]. For both the figures $\Omega_{\mathrm{m}0}=0.31$.
Left figure of Fig. 5 shows the matter power spectrum for MSE potential (4.1) for $n=6$ and $10$ along with the $\Lambda$CDM case.We can see that the power spectrum is getting suppressed as we decrease the value of $n$ and the linear perturbation level the MSE potential gives different predictions than the $\Lambda$CDM model. This nature is also represented in the right figure of Fig. 5 where the evolution of $f\sigma_{8}(z)$ has been shown. We can see clear difference in the evolution of $f\sigma_{8}(z)$ for the MSE potential than the standard $\Lambda$CDM model. For the other three potential (Eqs. (3.1), (5.12) and (5.13)) we don’t see any difference in the behaviour of power spectrum and $f\sigma_{8}(z)$ which clearly tells us that the difference between the MSE potential and $\Lambda$CDM does not come from the EDE solution but rather it comes from the late time behaviour of the scalar field. For SE (3.1), axion-like (5.12) and power law (5.13) potentials we have considered a CC along with the potential itself which is responsible foe the late time behavior. But, for the MSE potential (4.1) there is no need of additional CC term and viable late time dynamics is governed by the potential itself. Also, the late time behaviour of the scalar field in the MSE potential has tracker nature. This difference in the late time behaviour in MSE potential and $\Lambda$CDM model is responsible for the difference in the power spectrum and the evolution of $f\sigma_{8}(z)$ in the two cases. On the other hand, for the other three EDE potentials there is no difference in the late time behaviour with the $\Lambda$CDM model and thereby we do not observe any difference in the power spectrum and the evolution of $f\sigma_{8}(z)$.
@@
$\big{<}\delta(t,\vec{k})\delta(t,\vec{k}^{\prime})\delta(t,\vec{k}^{\prime\prime})\big{>}=\delta^{(3)}(\vec{k}+\vec{k}^{\prime}+\vec{k}^{\prime\prime})\mathcal{B}(t,k,k^{\prime}),~{}~{}~{}~{}~{}$ (6.11)
@@
$\delta(a,\vec{k})$ $=$ $\delta_{1}(a,\vec{k})+\frac{1}{2}\delta_{2}(a,\vec{k})$ (6.12)
@@
$\mathcal{B}(t,k,k^{\prime})$ $=$ $2F_{2}(\vec{k},\vec{k}^{\prime})\mathcal{P}(t,k)\mathcal{P}(t,k^{\prime})+\rm cyc\,.$ (6.13)
where, $\delta_{1}$ and $\delta_{2}$ are the first and second order density contrast. $F_{2}(\vec{k},\vec{k}^{\prime})$ denotes the second-order kernel which can be represented in terms of the growing ($D_{+}(a)$) and decaying ($D_{-}(a)$) mode as
$F_{2}(a,\vec{k}_{1},\vec{k}_{2})$ $=$ $\int_{a_{\mathrm{m}}}^{a}\mathrm{d}\tilde{a}^{\prime}~{}\frac{\mathcal{D}(a,a^{\prime})\mathcal{K}(a^{\prime},\vec{k}_{1},\vec{k}_{2})}{2a^{\prime 4}H^{2}(a^{\prime})W_{\mathrm{r}}(a^{\prime})}\,,$ (6.14)
$\mathcal{D}(a,a^{\prime})$ $=$ $\frac{D_{+}^{2}(a^{\prime})}{D_{+}^{2}(a)}\Big{(}D_{-}(a)D_{+}(a^{\prime})-D_{+}(a)D_{-}(a^{\prime})\Big{)}\,,$ (6.15)
where, $a_{\mathrm{m}}$ is an initial scale factor which can be fixed during the matter dominated era, $\mathcal{K}(a^{\prime},\vec{k}_{1},\vec{k}_{2})$ is the symmetrized kernel which related the the Fourier transform of the inhomogeneous part of the evolution equation of the second order density contrast [84] and $W_{\mathrm{r}}(a^{\prime})$ is the wronskian given by [84]
$W_{\mathrm{r}}(a)$ $=$ $D_{+}(a)\frac{\mathrm{d}D_{-}(a)}{\mathrm{d}a}-D_{-}(a)\frac{\mathrm{d}D_{+}(a)}{\mathrm{d}a}$ (6.16)

### v3

@@ 4.2 Background dynamics
Figure 3: Numerically evolved MSE potential (18) for $V_{0}=5\times 10^{10}\rho_{\rm c0}$, $n=6$, $\phi_{0}=0.7$ and $\Omega_{\mathrm{m}0}=0.3$. Initial conditions are $\phi_{i}=\phi_{0}-0.1\phi_{0}$ and $\phi_{i}^{\prime}=0.1$.
@@
Figure 4: Evolution of th background quantities for the MSE potential (18) for $V_{0}=5\times 10^{10}\rho_{\rm c0}$, $n=6$, $\phi_{0}=0.7$ and $\Omega_{\mathrm{m}0}=0.3$. For the upper and bottom left figures the initial conditions are $\phi_{i}=\phi_{0}-0.1\phi_{0}$ and $\phi_{i}^{\prime}=0.1$. For the bottom right figure the lines are for different initial conditions as mentioned in the figure.
@@
As we have already discussed that for $\phi/M_{\textrm{Pl}}\ll\phi_{0}^{1/n}$, i.e., at higher redshifts (early times), the MSE potential (18) reduces to the SE potential (15). So, the cosmological dynamics of the scalar field will be similar to the SE potential at higher redshifts or early times. Also, when $\phi/M_{\textrm{Pl}}\sim\phi_{0}^{1/n}$ the MSE potential is still very steep which can be understood from Fig. 3 which shows the numerical evolution of the MSE potential for $n=6$ and $\phi_{0}=0.7$. For the chosen $\phi_{0}$ we have $\phi_{0}^{1/n}\approx 0.94$ where the potential is very steep. Note that the potential has been shown in log scale in the Fig. 3. So, around $\phi_{0}$ we will have the effect of the steep potential which can lead to multiple frozen epochs in the scalar field dynamics as discussed in the previous section (Sec. 3) and has been shown in the upper figure of Fig. 1. For $\phi/M_{\textrm{Pl}}>\phi_{0}^{1/n}$, from Eq. (21), we see that
$V(\phi)\approx V_{0}\mathrm{e}^{-\gamma}\mathrm{e}^{\gamma\phi_{0}M_{\textrm{Pl}}^{n}/\phi^{n}}\,.$ (26)
@@
The above equation is very crucial to understand the late time dynamics of the scalar field. The MSE potential (18) behaves like the potential (26) before becoming a flat potential i.e., during the recent past. The potential (26) is a generalization of the potential $V(\phi)\sim\mathrm{e}^{M_{\textrm{Pl}}/\phi}$ which gives rise to tracker solution [8, 9]. So, the potential (26) would also give rise to a tracker solution. In tracker behaviour the scalar field dynamics does not exactly follow the background energy density like in scaling behaviour but decays a little slower than the background. Because of this slower decay process the scalar field energy density eventually takes over the background unlike the scaling behaviour in which the scalar field energy density follows the background forever. This tracker behaviour in the MSE potential (18) brings a slight difference in the scalar field dynamics compared to the SE potential (15). Now when $\phi/M_{\textrm{Pl}}\gg\phi_{0}^{1/n}$ the MSE potential (18) behaves like a CC which makes the scalar field to exit from the tracker behaviour and the scalar field EoS $w_{\phi}$ becomes very close to $-1$. So, the MSE potential not only can give rise to viable cosmology but also give interesting dynamics along with relating two different energy scales and thereby we do not need to invoke any CC in the theory. The steep nature of the potential gives multiple frozen epochs in the evolution of the scalar field energy density which can be very useful to achieve viable EDE solutions which we discuss in the next section (Sec. 5). In Fig. 4 we have shown the cosmological evolution of different cosmological parameters for the MSE potential (18).In the top right figure of Fig. 4 we see that we have two frozen epochs in the evolution of the scalar field energy density which is followed by a tracker behaviour. The bottom figure of Fig. 4 shows that the scalar field EoS $w_{\phi}$ is very close to $-1$ at $z=0$. The top left figure of the Fig. 4 shows the evolution of the density parameters. The bottom right figure of Fig. 4 shows the stability of solutions in the $\Omega_{\mathrm{m}}-\Omega_{\phi}$ plane. The three lines in the figure represent three different initial conditions and all the line moving towards a common future stable de Sitter solution, $\Omega_{\mathrm{m}}=0$ and $\Omega_{\phi}=1$. $\Omega_{\mathrm{m}}=\Omega_{\phi}=0$ represents radiation dominated era while $\Omega_{\mathrm{m}}=1$ and $\Omega_{\phi}=0$ represents matter dominated era. Both are either unstable or saddle points.
@@ 5 Early Dark Energy
$\theta_{\rm s}=\frac{r_{\rm s}(z_{\star})}{(1+z_{\star})D_{\rm A}(z_{\star})}\,,$ (27)
@@
$r_{\rm s}(z)$ $=$ $\int_{z}^{\infty}\frac{c_{s}(z)}{H(z)}{\rm d}z=\frac{1}{H_{\star}}\int_{z}^{\infty}\frac{c_{s}(z)}{H(z)/H_{\star}}{\rm d}z\,,~{}~{}~{}$ (28)
$D_{\rm A}(z)$ $=$ $\frac{c}{1+z}\int_{0}^{z_{\star}}\frac{dz}{H(z)}=\frac{c}{H_{0}(1+z)}\int_{0}^{z}\frac{dz}{E(z)}\,,~{}~{}~{}~{}$ (29)
@@
$c_{\rm s}(z)$ $=$ $\frac{c}{\sqrt{3(1+\frac{R_{\rm b}}{1+z})}}\,,$ (30)
$R_{\rm b}$ $=$ $\frac{3\omega_{\rm b}}{4\omega_{\gamma}}\,,$ (31)
$\omega_{\mathrm{r}}$ $=$ $\left(1+\frac{7}{8}N_{\rm eff}\left(\frac{4}{11}\right)^{4/3}\right)\omega_{\gamma}\,,$ (32)
@@
$z_{\star}$ $=$ $1048\left(1+0.00124\omega_{\rm b}^{-0.738}\right)\left(1+g_{1}\omega_{\mathrm{m}}^{g_{2}}\right)$ (33)
$g_{1}$ $=$ $\frac{0.0783\omega_{\rm b}^{-0.238}}{1+39.5\omega_{\rm b}^{0.763}}$ (34)
$g_{2}$ $=$ $\frac{0.560}{1+21.1\omega_{\rm b}^{1.81}}\;.$ (35)
@@
Table 1: Value of $H_{0}$ in the unit of ${\rm km}\;{\rm s}^{-1}{\rm Mpc}^{-1}$ for different $n$ by maintaining $\Omega_{\mathrm{m}0}\approx 0.31$, $f_{\rm EDE}\approx 0.07$, $z_{\star}\approx 1091$ and $z_{\rm eq}=3300$ and the redshift of drag epoch $z_{\rm d}=1060$. To calculate $H_{0}$ we have added a CC with SE (15), axionlike (38) and power law (39) potentials. Here $\lambda$ signifies $\mu$, $\phi_{0}$ and $f$.
@@
Using Eqs. (27), (28) and (29) we can have [11]
$H_{0}$ $=$ $\sqrt{3}H_{\star}\theta_{\rm s}\frac{\int_{0}^{z_{\star}}\frac{{\rm d}z}{E(z)}}{\int_{z_{\star}}^{\infty}\frac{E(z_{\star})}{E(z)}(1+R)^{-1/2}{\rm d}z}$ (36)
@@
$H_{\star}$ $=$ $\omega_{\mathrm{r}}^{1/2}(1+z_{\rm ls})^{2}\Big{(}\frac{\omega_{\mathrm{m}}}{\omega_{\mathrm{r}}}\times\frac{1}{1+z_{\star}}+1\Big{)}^{1/2}\times 100{\rm km~{}sec^{-1}Mpc^{-1}}\,.$ (37)
Eq. (36) gives us an estimate of $H_{0}$. We have listed values of $H_{0}$ for different parameter values for the SE (15) and MSE (18) potentials in Tab. 1 along with the values of $H_{0}$ for the axion-like and power law potentials given by
$V(\phi)$ $=$ $V_{0}\left(1-\cos\left(\frac{\phi}{f_{\rm pl}}\right)\right)^{n}\,,$ (38)
$V(\phi)$ $=$ $V_{0}\left(\frac{\phi}{f_{\rm pl}}\right)^{2n}\,,$ (39)
respectively by keeping the EDE density parameter at the redshift $z_{\rm c}$, $f_{\rm EDE}$ fixed at $0.07$ as small amount of $f_{\rm EDE}(<0.1)$ is allowed by the observation [40, 41, 38, 43, 24]. In the above equations $f_{\rm pl}=fM_{\textrm{Pl}}$ where $f$ is a constant. One should note that to estimate $H_{0}$ for using the Eq. (36) we are using the bounds from CMB observations on the parameters $\omega_{\mathrm{r}}$, $\omega_{\gamma}$, $\omega_{\mathrm{m}}$, $\omega_{\rm b}$ and $\theta_{\rm s}$ as mentioned above. So, this is just an estimation. The proper analysis should be done by doing observational data analysis by considering different cosmological observations. From Tab. 1 we see that the MSE potential (18) can improve the value of $H_{0}$ compared to $\Lambda$CDM but SE potential (15) with a CC improves the value even more. This happens because of the presence of tracker behaviour in the MSE potentials during the late time while we have scaling behaviour in the SE potential during the late time. Due to tracker behaviour the present value of $w_{\phi}$ can be slightly higher than $-1$ which can reduce the effect of EDE. While in SE potential, as we are adding a CC, the present value of $w_{\phi}$ is $-1$ which can have maximum effect of EDE for $w_{\phi}\geq 0$. It is also should be noted from Tab. 1 that the EDE in SE potential can give us better result in terms of the value of $H_{0}$ compared to the axion-like potential. To get the proper estimation of $H_{0}$ we do observational data analysis in the next section.
@@ 6 Growth of structures
$\ddot{\delta}+2H\dot{\delta}-4\pi G\bar{\rho}_{m}\delta=0\,,$ (41)
where, $\bar{\rho}_{\mathrm{m}}$ is the unperturbed (background) matter energy density and $\delta=(\rho_{\mathrm{m}}-\bar{\rho}_{\mathrm{m}})/\bar{\rho}_{\mathrm{m}}$ is the density contrast. The solution of Eq. (41) can be written as the linear superposition of the growing $D_{+}$ and decaying $D_{-}$ modes,
$\delta(t,\vec{k})=c_{+}D_{+}(t)\delta(\vec{k},0)+c_{-}D_{-}(t)\delta(\vec{k},0)\,,$ (42)
@@
$f=\frac{\mathrm{d}\ln D_{+}}{\mathrm{d}\ln a}\,.$ (43)
@@
$\Big{<}\delta(t,\vec{k})\delta(t,\vec{k}^{\prime})\Big{>}=\delta^{(3)}(\vec{k}+\vec{k}^{\prime})\mathcal{P}(t,k)\,,$ (44)
@@
$\mathcal{P}(t,k)$ $\propto$ $A_{\rm H}^{2}\Big{|}D_{+}(\tilde{a})\Big{|}^{2}T^{2}(k)\left(\frac{k}{H_{0}}\right)^{n_{\mathrm{s}}}\,,$ (45)
@@
$\sigma_{\rm R}^{2}=\int_{0}^{\infty}\mathrm{d}k\frac{k^{2}}{2\pi^{2}}\mathcal{P}(t,k)\big{|}W_{\rm win}(kR)\big{|}^{2}\,,$ (46)
@@
$\delta(\vec{x};R)=\int\delta(\vec{x}^{\prime})W_{\rm win}(\vec{x}-\vec{x}^{\prime};R)\mathrm{d}^{3}x^{\prime}\,.$ (47)
@@
$W_{\rm win}(kR)=\frac{3}{(kR)^{3}}\Big{(}\sin(kR)-kR\cos(kR)\Big{)}\,,$ (48)
The smoothing scale at which $\sigma_{\rm R}\sim 1$ is the scale where the linear perturbation theory breaks down and nonlinear effects become important. In this regard $R=8h^{-1}\rm Mpc$ is a relevant scale. From Planck 2018 results we have $\sigma_{8}=0.811\pm 006$ [3] at $z=0$. Using the best fit value of $\sigma_{8}$ from Planck 2018 results we fix the normalization factor $A_{\rm H}$ of the power spectrum  (45). $\sigma_{8}(z)$ can be represented by the growing mode $D_{+}(z)$
$\sigma_{8}(z)=\sigma_{8}(0)\frac{D_{+}(z)}{D_{+}(0)}\,.$ (49)
@@
Figure 5: Left: Power spectrum is shown for the MSE potential (18) with $n=6$ (solid blue) and $n=10$ (dotted green) at $z=0$ along with the power spectrum for the $\Lambda$CDM (dashed red) model. Right: Evolution of $f\sigma_{8}(z)$, for the similar cases as the left figure, has been shown. The brown dots are the observational data of $f\sigma_{8}$ along with their $1\sigma$ error bars [83]. For both the figures $\Omega_{\mathrm{m}0}=0.31$.
Left figure of Fig. 5 shows the matter power spectrum for MSE potential (18) for $n=6$ and $10$ along with the $\Lambda$CDM case.We can see that the power spectrum is getting suppressed as we decrease the value of $n$ and the linear perturbation level the MSE potential gives different predictions than the $\Lambda$CDM model. This nature is also represented in the right figure of Fig. 5 where the evolution of $f\sigma_{8}(z)$ has been shown. We can see clear difference in the evolution of $f\sigma_{8}(z)$ for the MSE potential than the standard $\Lambda$CDM model. For the other three potential (Eqs. (15), (38) and (39)) we don’t see any difference in the behaviour of power spectrum and $f\sigma_{8}(z)$ which clearly tells us that the difference between the MSE potential and $\Lambda$CDM does not come from the EDE solution but rather it comes from the late time behaviour of the scalar field. For SE (15), axion-like (38) and power law (39) potentials we have considered a CC along with the potential itself which is responsible foe the late time behavior. But, for the MSE potential (18) there is no need of additional CC term and viable late time dynamics is governed by the potential itself. Also, the late time behaviour of the scalar field in the MSE potential has tracker nature. This difference in the late time behaviour in MSE potential and $\Lambda$CDM model is responsible for the difference in the power spectrum and the evolution of $f\sigma_{8}(z)$ in the two cases. On the other hand, for the other three EDE potentials there is no difference in the late time behaviour with the $\Lambda$CDM model and thereby we do not observe any difference in the power spectrum and the evolution of $f\sigma_{8}(z)$.
@@
$\big{<}\delta(t,\vec{k})\delta(t,\vec{k}^{\prime})\delta(t,\vec{k}^{\prime\prime})\big{>}=\delta^{(3)}(\vec{k}+\vec{k}^{\prime}+\vec{k}^{\prime\prime})\mathcal{B}(t,k,k^{\prime}),~{}~{}~{}~{}~{}$ (50)
@@
$\delta(a,\vec{k})$ $=$ $\delta_{1}(a,\vec{k})+\frac{1}{2}\delta_{2}(a,\vec{k})$ (51)
@@
$\mathcal{B}(t,k,k^{\prime})$ $=$ $2F_{2}(\vec{k},\vec{k}^{\prime})\mathcal{P}(t,k)\mathcal{P}(t,k^{\prime})+\rm cyc\,.$ (52)
where, $\delta_{1}$ and $\delta_{2}$ are the first and second order density contrasts. $F_{2}(\vec{k},\vec{k}^{\prime})$ denotes the second-order kernel which can be represented in terms of the growing ($D_{+}(a)$) and decaying ($D_{-}(a)$) modes as
$F_{2}(a,\vec{k}_{1},\vec{k}_{2})$ $=$ $\int_{a_{\mathrm{m}}}^{a}\mathrm{d}\tilde{a}^{\prime}~{}\frac{\mathcal{D}(a,a^{\prime})\mathcal{K}(a^{\prime},\vec{k}_{1},\vec{k}_{2})}{2a^{\prime 4}H^{2}(a^{\prime})W_{\mathrm{r}}(a^{\prime})}\,,$ (53)
$\mathcal{D}(a,a^{\prime})$ $=$ $\frac{D_{+}^{2}(a^{\prime})}{D_{+}^{2}(a)}\Big{(}D_{-}(a)D_{+}(a^{\prime})-D_{+}(a)D_{-}(a^{\prime})\Big{)}\,,$ (54)
where, $a_{\mathrm{m}}$ is an initial scale factor which can be fixed during the matter dominated era, $\mathcal{K}(a^{\prime},\vec{k}_{1},\vec{k}_{2})$ is the symmetrized kernel which is related the Fourier transform of the inhomogeneous part of the evolution equation of the second order density contrast [84] and $W_{\mathrm{r}}(a^{\prime})$ is the Wronskian given by [84]
$W_{\mathrm{r}}(a)$ $=$ $D_{+}(a)\frac{\mathrm{d}D_{-}(a)}{\mathrm{d}a}-D_{-}(a)\frac{\mathrm{d}D_{+}(a)}{\mathrm{d}a}$ (55)